OpenStudy (anonymous):

11. Find the volume of a square pyramid with base edges of 24 cm and a slant height of 37 cm.

OpenStudy (accessdenied):

We can find the volume using the equation \( V = \frac{1}{3}b^2 h \). We know that \( b = 24 \) and we can find the height by using the slant height, \( s = 37 \), and half the base, \( \frac{1}{2} b = 12 \) with the Pythagorean Theorem \( a^2 + b^2 = c^2 \).